OpenStudy (anonymous):

What is the product of 8the square root of 16 times d. • 2dthe square root of 3.? Simplify if possible. group like terms?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

this is the expression right? \[\large 8\sqrt{16d} * 2d\sqrt{3}\]

OpenStudy (anonymous):

I think it's 8d sqrt 16 • 2d sqrt 3

OpenStudy (anonymous):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

actually one sec

jimthompson5910 (jim_thompson5910):

Rearrange the terms on the outside of the roots and multiply (steps 2 and 3). Then combine the roots (step 4) and rearrange terms inside the root (step 5). From there you break up the root and simplify. \[\large 8\sqrt{16d} * 2d\sqrt{3}\] \[\large 8*2d\sqrt{16d}*\sqrt{3}\] \[\large 16d\sqrt{16d}*\sqrt{3}\] \[\large 16d\sqrt{16d*3}\] \[\large 16d\sqrt{16*3d}\] \[\large 16d\sqrt{16}*\sqrt{3d}\] \[\large 16d*4*\sqrt{3d}\] \[\large 64d\sqrt{3d}\]

OpenStudy (anonymous):

I get it now! Thank you! I could actually see what you did each step to get that.

jimthompson5910 (jim_thompson5910):

That's great, I'm glad it's making sense